# The Fourth Conference on System and Service Quality - QualITA, June 25-26, 2025, Catania, Italy

### v1

@@ 4 Outlook
QMetric provides a structured approach for diagnosing hybrid quantum-classical models beyond conventional performance metrics. It highlights key aspects such as training behavior, encoding robustness, and circuit design quality. Future developments will include migration to Qiskit’s Estimator v2, support for additional platforms like PennyLane, and expanded metric coverage for multi-class tasks and generative models.
@@ 5 Availability
All source code, examples, metric definitions, and plotting utilities are available at: https://gitlab.com/illesova.silvie.scholar/qmetric
The repository includes a Conda environment file for reproducibility.
@@ 6 Acknowledgments
Finanziato dall’Unione Europea - Next Generation EU, Missione 4 Componente 1 CUP D53D23017570001 e E53D23016370001 e dal Governo italiano (Ministero dell’Università e della Ricerca, PRIN 2022 PNRR) – cod. P2022SELA7: “RECHARGE: monitoRing, tEsting, and CHaracterization of performAnce Regressions” – Decreto Direttoriale n. 1205 del 28/7/2023. TR gratefully acknowledges the funding support by program "Excellence initiative—research university" for the AGH University in Krakow as well as the ARTIQ project: UMO-2021/01/2/ST6/00004 and ARTIQ/0004/2021.

### v2

@@ 5 Outlook
QMetric provides a structured approach for diagnosing hybrid quantum-classical models beyond conventional performance metrics. It highlights key aspects such as training behavior, encoding robustness, and circuit design quality. Future developments will include migration to Qiskit’s Estimator v2888https://quantum.cloud.ibm.com/docs/en/api/qiskit-ibm-runtime/estimator-v2, support for additional platforms like PennyLane, and expanded metric coverage for multi-class tasks and generative models.
@@ 6 Availability
All source code, examples, metric definitions, and plotting utilities are available at https://gitlab.com/illesova.silvie.scholar/qmetric
The repository includes a Conda environment file to reproduce the case study.
@@ 7 Acknowledgments
MB was supported by Italian Government (Ministero dell’Università e della Ricerca, PRIN 2022 PNRR)—cod. P2022SELA7: “RECHARGE: monitoRing, tEsting, and CHaracterization of performAnce Regressions”—D.D. n. 1205 del 28/7/2023. TR gratefully acknowledges the funding support by program "Excellence initiative—research university" for the AGH University in Krakow as well as the ARTIQ project: UMO-2021/01/2/ST6/00004 and ARTIQ/0004/2021.